Devin is 51 years younger than Jamie. 7 years ago Jamie’s age was 4 times devins age. How old is devin now?
Found 2 solutions by ikleyn, hoomanc:
Answer by ikleyn(52788)

.

Let x be the Devin's age 7 tears ago.


Then Jamie's age 7 years ago was 4x, according to the condition,

and we are given additionally that

    4x - x = 51.


Hence,

    3x = 51, 

     x = 51/3 = 17.


Thus 7 years ago Devin was 17 years old.


ANSWER.  Devin is 17+7 = 24 years old now.

Solved.

-----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.




Answer by hoomanc(14)


D= Devon in 2020
,J= Jamie in 2020
(D-7)= Devon in 2016
(D-7)= Devon in 2016
D+51=J
4*(D-7)=(J-7)
Try to subtract the two side of above equation together to remove J from right side.
D+51-(4D-28)=J-(J-7)
-3D+79=7 >>> 72=3D >>> D=24
And from first equation we have >>> J=D+51 >>> J=75